### v1

# Exploring selection biases in Fast Radio Burst (FRB) dispersion-galaxy cross-correlations with magnetohydrodynamical simulations

Abstract
The dispersion of fast radio bursts (FRBs) in conjunction with their redshifts can be used as powerful probes of the distribution of extragalactic plasma, and with a large enough sample, the free-electron–galaxy power spectrum $P_{eg}$ can be measured by cross-correlating FRB dispersions with galaxy positions. However, a precise measurement of $P_{eg}$ requires a careful investigation of the selection effects—the fact that the probability of both observing the FRB dispersion measure and obtaining a host galaxy redshift depends on observed properties of the FRB and its host. We use ray tracing simulations with IllustrisTNG300-1 to investigate the impact of expected observational selection effects on FRB dispersion–galaxy angular cross-correlations with a sample of 3000 FRBs at redshift range of $0.3\leq z\leq 0.4$ Our results show that cross-correlations with such an FRB sample are robust to properties of the FRB host galaxy: this includes DM contributions from the FRB host and optical followup selection effects biased against FRBs with dim galaxy hosts. We also find that such cross-correlations are robust to DM dependent and scattering selection effects specific to the CHIME/FRB survey. However, a DM dependent selection effect that cuts off the 10% most dispersed FRB at a fixed redshift shell can bias the amplitude of the cross-correlation signal by over 50% at angular scales of $\sim 0.1^{\circ}$ (corresponding to Mpc physical scales). Our findings highlight the importance of both measuring and accounting for selection effects present in existing FRB surveys as well as mitigating DM dependent selection effects in the design of upcoming FRB surveys aiming to use FRBs as probes for large-scale structure.
1 Introduction
Fast Radio Bursts (FRBs) are extragalactic millisecond-duration radio transients originating at cosmological redshifts. They are characterized by their large dispersive delays, parameterized via dispersion measures (DM), where the dispersion measure (DM) is directly proportional to the integrated free electron density along the line of sight. Since the majority of the Universe’s baryonic matter is in the form of a diffuse plasma outside of galaxies, FRB dispersion is an excellent tracer of baryons. Unlike other tracers of baryonic matter, such as X-ray observations and the thermal Sunyaev-Zeldovich effect, which have a temperature-dependent signal, FRB DMs probe hot and cold ionized gas equally. As such, FRBs have emerged as promising cosmological probes that can help solve a range of astrophysical problems that are sensitive to the distribution of baryonic matter (Lorimer et al., 2007; Thornton et al., 2013; McQuinn, 2014; Macquart et al., 2020), from improving galaxy-cluster kinetic SZ measurements (Madhavacheril et al., 2019) to constraining models of astrophysical feedback (Medlock et al., 2024; Sharma et al., 2025).
While to date thousands of FRBs have been detected (CHIME/FRB Collaboration et al., 2021a), only $\sim 100$ have been localized with the angular resolution necessary to identify their host galaxies and obtain redshifts (Hallinan et al., 2019; Bhandari et al., 2020; FRB Collaboration et al., 2025a). Nevertheless, enough FRB redshifts have been measured to enable studies of the extragalactic electron density distribution through the FRB line of sight via “one-point” statistics of DM distributions. This typically requires partitioning the total dispersion measure of an FRB into its constituent parts (e.g. a contribution from the FRB host galaxy, a cosmic contribution from the CGMs intervening halos and the IGM, and a contribution from the Milky Way) (Connor et al., 2024; Sharma et al., 2025) that assume a parameterization for each component based on hydrodynamical simulations. Parameters of interest are then used to probe a given astrophysical process; for instance, the so called “F-parameter” originally used by Macquart et al. (2020) and then again by Connor et al. (2024) to constrain the fraction of all baryons in collapsed halos, parametrizes the variance of the cosmic contribution of the DM as $\sigma_{DM,\rm{cosmic}}=Fz^{-1/2}$, while Sharma et al. (2025) proposes parameterizing the variance as a redshift dependent log-normal distribution which can in turn be used to constrain the strength of different astrophysical feedback processes.
In general there are two main disadvantages to using one-point statistics of the DM. The first is that the precision of the DM-budgeting procedure is limited by the accuracy of the (likely overly simplistic) underlying parametrization of each DM component, which can be motivated but not verified by simulations. The second is sensitivity to selection effects—the fact that the probability of detecting an FRB depends on its observed properties—which can bias the baryon distribution as measured from FRB DMs. For instance, both the ASKAP and CHIME/FRB search pipelines are insensitive to over 50% of the expected FRB population at DM$>$1000pc/cm-3 (Shannon et al., 2018; Shin et al., 2023; Merryfield et al., 2023). If, as a simple example, this preferentially removes FRB detections that intersect halos with a dense electron density, a measurement of the distribution of baryons in collapsed halos using the one-point statistic would be an underestimate.
[…]
In light of thousands to tens of thousands of FRB redshifts on the horizon enabling increasingly more precise measurements of the extragalactic electron distribution, this work aims to investigate the extent to which selection biases in upcoming FRB surveys can be expected to limit the precision of the dispersion measure-galaxy cross-correlation using IllustrisTNG. IllustrisTNG is a suite of large-scale gravomagnetohydrodynamical simulations varying in resolution, physics complexity, and the size of its cubic simulation box, enabling the study of cosmological phenomena across different physical scales (Vogelsberger et al., 2014; Springel et al., 2018; Pillepich et al., 2018; Nelson et al., 2019). We conduct our simulation with IllustrisTNG’s largest simulation box size of $300\,\mathrm{Mpc}$ containing the largest number of galaxies, TNG300-1.
In this work, we demonstrate through our ray tracing simulations which selection biases cross-correlations are the most sensitive to, along with the rough magnitudes of the effects. In Section 2 we describe the details of our cross-power spectrum estimation and ray tracing simulations, and in Section 3 we present and discuss the results of our different simulated selection biases. We conclude in Section 4.
2 Preliminaries and Methods
2.1 The DM-galaxy cross-correlation power spectrum
The dispersion measure of an FRB observed at a comoving distance $\chi$ and sky location $\vec{\theta}$ is given by
[…]
where $n_{e}$ denotes the comoving electron number density and $z$ is the redshift. Therefore, the average electron number density $\bar{n}_{e}(z)$ gives us $\overline{\mathrm{DM}}(z)$, the expected DM of an FRB at redshift $z$; this is the well-known Macquart relation (Macquart et al., 2020). Given a catalog of FRBs, we can then define the DM overdensity field as
$d(\vec{\theta},z)=\mathrm{DM}(\vec{\theta},z)-\overline{\mathrm{DM}}(z),$ (2)
The DM overdensity is thus a measurement of plasma overdensities integrated along the FRB line of sight. Indeed, we can similarly define the galaxy overdensity field as
$g(\vec{\theta},z)=\frac{n_{g}(\vec{\theta},z)-\overline{n}_{g}(\vec{\theta},z)}{\overline{n}_{g}(\vec{\theta},z)},$ (3)
where $n_{g}(\vec{\theta},z)$ is the observed galaxy number density at $(\vec{\theta},z)$, and $\overline{n}_{g}(\vec{\theta},z)$ is the expected number density in the absence of clustering, which can be a function of sky position if there are non-uniform survey selection effects.
Given the DM and galaxy count overdensity fields, the quantity we are interested in measuring is the angular cross-power spectrum, defined as
$\delta_{ll^{\prime}}\delta_{mm^{\prime}}C^{Dg}_{\ell}(z_{host},z_{g})=\langle\tilde{d}^{*}_{lm}(z_{host})\tilde{g}_{l^{\prime}m^{\prime}}(z_{g})\rangle,$ (4)
[…]
$C^{Dg}_{\ell}=n_{e,0}\frac{1+z_{g}}{\chi_{g}^{2}}P_{eg}(k=\ell/\chi_{g},z_{g}).$ (5)
where $P_{eg}$ in turn probes the distribution of baryonic matter around galaxies.
2.2 Power Spectrum Estimation
In practice, we discretize the survey volume, and quantities $\mathrm{DM}$, $n_{g}$, etc. are evaluated by binning the FRBs and galaxies into angular sky pixels and tomographic redshift bins. Note that an FRB survey only yields sparse measurements of the DM field at the sky locations of the survey FRBs. That is, $\mathrm{DM}(\vec{\theta},z)$ only has a measured value in pixels that contains at least one FRB; an average is taken if there is more than one FRB in a pixel.
Our mock catalogs examine the sky in square patches of approximately $6^{\circ}\times 6^{\circ}$, a choice required by the geometry of the simulation box (see Appendix B for a detailed discussion). Therefore, we can use the flat sky approximation, in which the SHT reduces to a 2D discrete Fourier transform (DFT):
$\tilde{f}(\vec{\ell})=\frac{A}{N^{2}}\sum_{x=0}^{N-1}\sum_{y=0}^{N-1}f(\vec{\theta}_{x,y})\,\exp(-i\,\vec{\ell}\cdot\vec{\theta})$ (6)
for some field $f(\vec{\theta})$, where $A$ is the angular area of the sky patch and $N$ is the number of pixels on a side. The flat sky angular power spectrum is then
$\delta_{\vec{\ell}\vec{\ell}^{\prime}}C_{l}^{Dg}=\frac{1}{A}\langle\tilde{d}^{*}_{\vec{\ell}}\,\tilde{g}_{\vec{\ell}^{\prime}}\rangle.$ (7)
Naively, the power spectrum above can be estimated by averaging over the independent Fourier modes [^1] For any real field $f$, the reality condition imposes that $\tilde{f}_{\vec{\ell}}=\tilde{f}^{*}_{-\vec{\ell}}$. $\vec{\ell}$ that satisfy $|\vec{\ell}|=\ell$:
$\hat{C}_{\ell}^{Dg}=\frac{1}{AN_{\ell}}\sum_{|\vec{\ell}|=\ell}\tilde{d}^{*}_{\vec{\ell}}\,\tilde{g}_{\vec{\ell}},$ (8)
or, if we bin into bandpowers $\lambda$,
$\hat{C}_{\lambda}^{Dg}=\frac{1}{AN_{\lambda}}\sum_{|\vec{\ell}|\in\lambda}\tilde{d}_{\vec{\ell}}\,\tilde{g}^{*}_{\vec{\ell}},$ (9)
where $N_{\lambda}$ is the total number of independent modes belonging to bandpower $\lambda$.
However, in any realistic survey, we will not have an FRB in every sky pixel, and thus the DM measurement is incomplete. This is equivalent to multiplying the DM field with a boolean window function $w^{d}_{\vec{\theta}}$ corresponding to whether or not at least one FRB was detected at $\vec{\theta}$. Similarly, the galaxy field can have its own survey window $m^{g}_{\vec{\theta}}$. With incomplete surveys, the naive estimator defined in Equation 8 can only operate on the masked fields, namely
$\hat{C}_{\lambda}^{Dg}=\frac{1}{AN_{\lambda}}\sum_{|\vec{\ell}|\in\lambda}(\tilde{d}^{m}_{\vec{\ell}})^{*}\,\tilde{g}^{m}_{\vec{\ell}},$ (10)
where $d^{m}_{\vec{\theta}}=m^{d}_{\vec{\theta}}\,d_{\vec{\theta}}$ and $g^{m}_{\vec{\theta}}=w^{g}_{\vec{\theta}}\,g_{\vec{\theta}}$ are the masked DM and galaxy overdensity fields. In this case, the estimated power spectrum usually suffers from mixing of power across different angular scales, i.e.,
$\hat{C}^{Dg}_{\lambda}=\sum_{\lambda^{\prime}}M_{\lambda\lambda^{\prime}}\overline{C}^{Dg}_{\lambda^{\prime}},$ (11)
where $\overline{C}^{Dg}_{\lambda^{\prime}}$ is the true underlying angular power spectrum, and $M_{\lambda\lambda^{\prime}}$ is the power mixing matrix which depends on the survey windows $m^{d}$ and $m^{g}$.
To obtain an unbiased estimate for the cross-power spectrum, we can use the optimal quadratic estimator (OQE) derived in Tegmark (1997):
$C_{\lambda}^{Dg}=\sum_{\lambda^{\prime}}(F_{\lambda\lambda^{\prime}})^{-1}\,\frac{(\mathbf{g}^{m})^{\dagger}\mathbf{G}^{-1}\mathbf{C}_{,\lambda^{\prime}}\mathbf{D}^{-1}\mathbf{d}^{m}}{2}.$ (12)
Here, $\mathbf{g}^{m}$ and $\mathbf{d}^{m}$ are the masked real space galaxy and DM overdensity fields flattened into vectors and $\mathbf{D}=\langle\mathbf{d}^{m}(\mathbf{d}^{m})^{\dagger}\rangle$, $\mathbf{G}=\langle\mathbf{g}^{m}(\mathbf{g}^{m})^{\dagger}\rangle$, and $\mathbf{C}=\langle\mathbf{d}\mathbf{g}^{\dagger}\rangle$ are the DM, galaxy, and DM-galaxy (cross-)covariance matrices, respectively. $\mathbf{C}_{,\lambda^{\prime}}=\partial\mathbf{C}/\partial C_{\lambda^{\prime}}^{Dg}$ is the derivative cross-covariance matrix, which can be written in terms of the DFT operators on the masked fields and whose form we derive in Appendix A. $F_{\lambda\lambda^{\prime}}$ is the Fisher information matrix, which here is
$F_{\lambda\lambda^{\prime}}=\frac{1}{2}\mathrm{Tr}[\mathbf{C}_{,\lambda}\mathbf{G}^{-1}\mathbf{C}_{,\lambda^{\prime}}\mathbf{D}^{-1}].$ (13)
In a realistic survey, the full covariances $\mathbf{D}$ and $\mathbf{G}$ can be difficult to estimate. Even if they can be obtained, the computational cost for Equation 13 is often too high. Works such as Alonso et al. (2019) have shown that simplifications by assuming the covariances $\mathbf{D}^{-1}$ and $\mathbf{G}^{-1}$ are diagonal allow Equations 12 and 13 to be computed efficiently while keeping the estimator close to optimal in most cases. One simple way to estimate $\mathbf{D}^{-1}$ and $\mathbf{G}^{-1}$ is to have some estimated variance $\sigma^{-2}$ along the diagonal for pixels with at least one FRB, and 00 for pixels with no FRBs, corresponding to a limit of infinite noise (i.e. the DM is not measured). However, any constant that multiplies $\mathbf{D}^{-1}$ or $\mathbf{G}^{-1}$ will cancel out between the inverse Fisher matrix and the numerator in Equation 12. Therefore, we can simply take the diagonal of $\mathbf{D}^{-1}$ and $\mathbf{G}^{-1}$ to be the survey mask:
$\mathbf{D}^{-1}=\text{diag}(\mathbf{m}^{d})$ (14)
$\mathbf{G}^{-1}=\text{diag}(\mathbf{m}^{g}),$ (15)
where $\mathbf{m}^{d}$ and $\mathbf{m}^{g}$ are the FRB survey mask and galaxy survey mask flattened into 1-D vectors, respectively. As a sanity check, we show in Appendix A that this estimator reduces to the naive estimator in Equation 9 in the limit of complete sky coverage.
Finally, the error on the OQE, assuming the fields are Gaussian and that there is no window function present, can be shown to be (Tegmark, 1997)
$\langle(C_{\lambda}^{Dg}-\overline{C}_{\lambda}^{Dg})^{2}\rangle=(\Delta C_{\lambda})^{2}=\frac{\overline{C}^{DD}_{\lambda}\overline{C}^{gg}_{\lambda}+\left(\overline{C}^{Dg}_{\lambda}\right)^{2}}{N_{\lambda}},$ (16)
where $C^{DD}_{\lambda}$ and $C^{gg}_{\lambda}$ are the auto power spectrum for the DM and galaxy overdensity field, respectively. However, as we discuss further in Section 2.4, the Gaussianity assumption on which this equation relies is unlikely to be valid on small scales due to non-linear structure growth. We therefore estimate the errors of our cross-power spectrum empirically, by computing it over multiple realizations in many independent sky patches (see Section 2.3).
2.3 Ray Tracing with IllustrisTNG
The IllustrisTNG project is a suite of large-scale, cosmological magnetohydrodynamical simulations that model galaxy formation and evolution across cosmic time. IllustrisTNG consists of three volumes with box sizes 50, 100, and 300 cMpc—where c denotes comoving distance–enabling cosmological studies focusing on different scales; each box size is run at three different resolution levels. For our study, we choose TNG300-1, the highest resolution run of the largest simulation box in order to simulate FRBs found at cosmological distances. Each simulation comes in 100 discrete “snapshots” containing field data [^2] See https://www.tng-project.org/data/docs/specifications. of all particles in the simulation box, of which 80 are “mini” snapshots that contain a subset of the field particle information as the “full” snapshots. We access the data via the online IllustrisTNG JupyterLab workspace [^3] https://www.tng-project.org/data/lab.
The size of such a cosmological simulation is extremely large. The simulation box of TNG300-1 consists of $2500^{3}$ particles, with a total size of 4.1 TB per full snapshot. Directly ray tracing through the simulation data would require a linear search of all particles in the simulation box along the given sightline, a computationally infeasible task for even a single FRB. Previous works (Zhang et al., 2021; Walker et al., 2024) have compressed the data by choosing a single direction along which to ray trace and pre-computing the DM along those sightlines. Because we need to be able to place FRBs in arbitrary directions, we instead compress the data by computing the electron density field of each snapshot on a grid. To do so, we sort the particles of a given snapshot into cubic bins. The electron number count of each particle is then computed as
$N_{e}=\eta_{e}\,X_{H}\frac{\rho}{m_{p}}V$ (17)
where $\eta_{e}$ is the electron abundance with respect to the hydrogen number density, $X_{H}$ is the total hydrogen abundance, $\rho$ is the comoving mass density, $m_{p}$ is the proton mass, and $V$ is the volume of the particle. The comoving electron density of each bin is obtained by dividing its total electron number count by its volume. We then repeat this for each snapshot. Note that while $X_{e}$ and $\rho$ are available in all snapshots, $X_{H}$ is only available in the full snapshots, and we take $X_{H}=0.76$ for the mini-snapshots. We have checked and found no discontinuities in the line-of-sight electron density as a result of this approximation.
[…]
Another difficulty in ray tracing in IllustrisTNG comes from the fact that FRBs are generally observed to come from cosmological distances ($z\gtrsim 0.1$) (Macquart et al., 2020) that exceed the $300\,\mathrm{cMpc}$ size of the simulation box. Since the simulation box has periodic boundary conditions, given observer and FRB positions, computing the FRB DM can be done by stacking simulation snapshots. However, this raises the possibility of a sightline (or nearby sightlines within the same sky region) intersecting the same part of the simulation box multiple times. Previous works have dealt with this issue by either changing the angle of the line of sight segment or performing linear coordinate transformations (e.g. rotating the box) when the ray crosses the box boundary (Zhang et al., 2021; Walker et al., 2024). While this solution is sufficient for studies of DM distributions, the cross-correlation is a statistic that explicitly describes spatial correlations, and therefore requires the preservation of relative positions of particles within the light cone. We resolve this problem geometrically, computing a-priori the “good” regions of sky, where no sightline intersects with itself or its neighboring pixels for a given pixel resolution.
The larger the distance to a given FRB, the more times the simulation box must be stacked and therefore the more likely it is that a given line of sight will intersect with the same structure more than once. Equivalently, the larger the FRB redshift, the smaller the allowed sky regions become, therefore requiring a compromise between these quantities. We leave the details of these computations to Appendix B. The result is summarized in Figure 1, which shows the allowed sky regions for our choice of $z_{\max}=0.4$ and pixel resolution $0.0008\,\mathrm{rad}$, along with the $48$ $0.1\,\mathrm{rad}\times 0.1\,\mathrm{rad}$ square sky patches within the allowed regions. There are $15625125^{2}=$ 15625 pixels per sky region.
Figure 1: The $48$ square sky regions (shaded gray) used in this study (out to $z_{\max}=0.4$) for independent realizations of the DM and galaxy field. The regions are chosen from the “good” regions of the sky (white) that avoid spurious geometrical effects from the simulation’s periodic boundary conditions.
Finally, our ray tracing procedure can be described as follows. Given any observer and FRB position, we:
  1. 1.
    Compute the total path of the ray through the box by evolving the ray from the observer to the FRB.
  2. 2.
    For each cubic bin the ray intersects, retrieve its electron density from the closest snapshot. The closest snapshot is defined as the snapshot with the smallest comoving distance between the center of the ray within the bin and the redshift at which the snapshot was taken.
  3. 3.
    Compute and save the cumulative $\mathrm{DM}(\chi)$ with Equation 1 using trapezoidal Riemann integration:
    $\mathrm{DM}(\chi)=\sum_{\chi_{i}\leq\chi}n_{e}(\chi_{i})\bigg{(}1+z(\chi_{i})\bigg{)}\Delta\chi_{i}.$ (18)
    where $\Delta\chi_{i}$ is the comoving length of the ray in the $i$th cubic bin, accounting for “partial intersections”.
By computing the cumulative line-of-sight DM, we only have to ray trace once for each pixel of each sky region. Then, we can compute the DM of an FRB placed within a sky region at an arbitrary $z<0.4$ via interpolation with the radial coordinate. We choose the same observer position for all sky regions, which is the lowest density region of the simulation box as our origin, since the binning of the simulation box wipes out nearby structures. Effectively, this is an explicit removal of any galactic contribution to the DM, corresponding to the assumption that the Milky Way DM is known and can be subtracted in a real survey.
Finally, the DM-galaxy cross-correlation requires a catalog of foreground galaxies. Because there are far fewer galaxies than simulation particles, it is not necessary to bin the galaxy map as we did for electrons. IllustrisTNG provides galaxies in the form of a “group catalog” at each snapshots, containing a list of halos and subhalos. Galaxies are taken to be subhalos flagged as galaxies (“SubhaloFlag=0”) with nontrivial stellar content ($M_{g}<0$). We build the light cone of galaxies in a similar fashion as our ray tracing procedure, stacking simulation boxes and evolving the cone in tomographic slices corresponding to each snapshot. We identify the redshift range each snapshot, computed as the halfway point (in comoving distance) to neighboring snapshots. Then, for each snapshot $i$, we create a galaxy catalog corresponding to that slice by taking all galaxies with $\chi(z_{\min,i})\leq|\vec{x}_{g}-\vec{x}_{\text{obs}}|<\chi(z_{\max,i})$, where $\vec{x}_{g}$ and $\vec{x}_{\text{obs}}$ are the positions of the galaxy and observer, respectively. We save the galaxies’ center-of-mass coordinates, along with other properties retrieved from IllustrisTNG (stellar mass, star formation rate (SFR), and $g$-band absolute magnitude), from which we straightforwardly compute the projected sky position and apparent magnitude. The end product is a catalog of galaxies within the cone of each sky region up to a maximum redshift.
Combined, these tools enable generic studies of realistic FRB surveys. In particular, we now use this framework to study potential systematics in the DM-galaxy cross-correlation in realistic FRB and galaxy surveys.
2.4 Fiducial cross-power spectrum
Using the methods described in the previous two sections, we populate the simulation volume with FRBs, use our ray-tracing methods to determine their DMs, and measure $C_{\ell}^{Dg}$ with our cross-power spectrum estimator. We now describe the parameters of our fiducial population model, without selection effects, and compare the resulting power-spectrum to theoretical expectations.
In all our experiments, we simulate $N=3000$ FRBs within a redshift slice $0.3<z_{\text{host}}\leq 0.4$ for each sky region shown in Figure 1. We draw our FRBs randomly from galaxies weighted by the galaxy’s SFR under the simple model that FRBs trace the star formation rate:
$W_{\text{host}}=\frac{\mathrm{SFR}_{\text{host}}}{1+z_{\text{host}}}$ (19)
where the $1+z_{\text{host}}$ arises from the cosmological time dilation of the rate. We then cross-correlate the DMs of our FRBs with foreground galaxies in a $0.2<z_{g}\leq 0.3$ redshift slice.
[…]
Figure 2: Comparison of $P_{eg}(z_{g})$ computed directly from the electron density and galaxy overdensities at redshift $z_{g}=0.2$ (black) to the electron galaxy power spectrum calculated from $C_{\ell}^{Dg}$ using Equation 5 (red). The top panel shows the ratio of the two power spectra, with the error bars calculated from the variance of different realizations of $C_{\ell}^{Dg}$. This demonstrates the validity of the theory that relates the observable dispersion angular power spectrum to the underlying three dimensional power spectrum of free electrons.
We estimate the uncertainties on the cross-correlation empirically; each of the $48$ sky regions represents an independent realization of the matter field, and we have $n_{\text{trial}}=5$ realizations of the FRB survey (via independent draws from galaxies) for each region, for a total of $240$ realizations of $C_{\ell}^{Dg}$. All figures show the mean and standard deviation $C_{\ell}^{Dg}$ from these $240$ total realizations unless otherwise specified.
We also note that the error bars (standard deviation over all $C^{Dg}_{\ell}$) placed on our power spectra are the error bars corresponding to a survey of 3000 FRBs, not the uncertainties on the reported mean simulated power spectrum measurement (the standard error) which are a factor of $\sqrt{48}$ smaller. The factor of $\sqrt{48}$ can be understood as follows: the total contribution to the cross correlation noise in an ideal survey of FRBs with redshifts (i.e. the Maquart relation is removed) can be decomposed into (a) sample/cosmic variance (b) FRB host galaxy/local environment noise and (c) foreground galaxy shot noise. In the case of our simulation where all foreground galaxies are used in the cross correlation, the galaxy shot noise can be ignored, leaving contributions (a) and (b). While each of the 5 trials randomizing the choice of background FRBs reduce the contribution of (b), the large-scale structure sample variance is the same for each trial within a given sky region, meaning we have only 48 effective realizations of (a).
Finally, we note that forecasts (e.g. Madhavacheril et al. (2019)) often make two central assumptions to obtain theoretical error bars (as given by Equation 16) on $C_{\ell}^{Dg}$: the first is that no window function is present, and the second is that the DM and galaxy fields are gaussian. In order to isolate the effect of the second–the non-gaussianity of the cosmic strucutre–we compare the actual standard deviation of the power spectrum for a sample of 15,625 FRBs (one FRB per pixel) to the expectation for Gaussian fields in Figure 3. We find that the Gaussian expectation underestimates the true error particularly at large $\ell$, which is a natural consequence of the fact that matter over densities become increasingly non-Gaussian at small scales due to non-linear structure growth. As such, future forecasts of this measurement should adjust error bars accordingly—although in practice, with real data the errors are computed using mock or randomized galaxy and FRB catalogs (see e.g. Rafiei-Ravandi et al. (2021), Wang et al. (2025, in prep.).
Figure 3: Comparison of theoretical expectation of the error on $C_{\ell}^{DG}$ as predicted by Equation 16 in the case of a Gaussian field to the true standard deviation of our measured $C_{\ell}^{DG}$ over many realizations for FRBs cross correlated with galaxies. Here, we have FRBs drawn uniformly distributed over the sky (one FRB/pixel) to avoid the additional effect the window function has on the variance. The true variance of the data is increasingly larger than the Gaussian expectation at smaller scales.
Figure 4: Comparison of cross correlation measurement when FRBs are drawn uniformly over all pixels (red) to when FRBs are drawn based on SFR (Equation 19). That FRBs stocastically—rather than uniformly—sample the DM field induces no bias on the power spectrum.
3 Impact of Selection Effects
3.1 Insensitivity to FRB and host galaxy properties
Figure 5: Top: Apparent magnitude distribution for fiducial sample of FRBs in the redshift range ($0.3<z_{g}\leq 0.4$). Bottom: Host galaxy incompleteness does not significantly affect the cross-correlation, even for an aggressive $m_{g}$ cut.
Figure 6: Top: The DM distribution of our fiducial sample of FRBs and our sample after injecting a host DM contribution. Bottom: The cross-correlation with the fiducial model and after injecting a host DM. The ratio of the error after injecting a host DM to the error on the fiducial model is shown in the second plot from the top, and the ratio of the $C^{Dg}_{\ell}$s are given in the third plot from the top. While the host DM contribution contributes noise to $C_{\ell}^{Dg}$, it does not significantly bias the cross correlation measurement.
Figure 7: Top: The change in DM distribution after introducing a DM-dependent selection function as given by Equation 21 for $a=2$, which roughly scales the DM-dependent selection effect from the CHIME/FRB survey down to our FRB redshift range. Bottom: The cross-correlation of the DM field with the DM-dependent selection effect applied, compared to the default model as reference. The residuals are plotted in the middle panel. A smooth DM-dependent selection effect not significantly affect the cross-correlation signal.
Figure 8: Top: The change in DM distribution after introducing a DM-dependent selection function as given by Equation 21 for $a=5$ (solid blue) and a step-wise selection function that cuts off at 800 pc cm-3 (red). The DM-dependent selection function for $a=2$ shown in Figure 7 to represent CHIME’s selection effects is plotted with the dashed blue line for reference. Bottom: The cross-correlation of the DM field with the DM-dependent selection effect applied, compared to the default model. The residuals are plotted in the middle panel. The shaded boxes and error bars indicate the standard deviation for the CHIME-like and step-wise DM dependent selection functions, respectively. With a DM cut given by Equation 21 with $a=5$, the amplitude of the power spectra drops by over 20% beyond $\ell\sim 10^{3}$. For the abrupt DM cut, the amplitude of the power spectra drops by over 50% beyond $\ell\sim 10^{3}$.
Because, in principle, the signal in the cross-correlation arises purely from the electrons along the FRB line of sight that are clustered around foreground galaxies, it is convenient to consider two different kinds of selection effects. The first are selection effects on observational properties of the FRB and host galaxy that arise outside of the foreground galaxy redshift shell. The second are selection effects on observational properties of the FRB that arise within the galaxy redshift shell, which we will refer to as “propagation selection effects”. In this section, we consider the former and in section 3.2 we consider the latter.
3.1.1 FRB population model
We adopt a distribution of FRBs that trace the SFR of galaxies within a specified redshift shell in our fiducial model. As discussed in Section 2.1, this introduces a window function on the DM overdensity map that mixes power on different angular scales, which can be mitigated with the OQE. We test the effect of the physical placement of FRBs on the sky in Figure 4 by comparing our measurement of $C_{\ell}^{Dg}$ with the fiducial FRB sample (criteria listed in Section 2.4) to our measurement with an identical sample of FRBs except with FRBs uniformly drawn over the sky (one FRB per pixel). We find that the measurement of $C_{\ell}^{Dg}$ is consistent when drawing FRBs uniformly over the sky and when drawing them based on SFR.
3.1.2 Optical followup selection effects
Redshifts can only be obtained for FRBs if they are associated with a galaxy that is bright enough to be observable. FRBs originating from very dim or even “unseen hosts” (see eg. Marnoch et al., 2023) then will be selected against in any real sample of localized FRBs with redshift measurements.
As done in Section 3.1.1, we test the hypothesis that the cross correlation is robust to selection effects that bias against detecting FRBs in dim hosts directly with our simulation. Although the current magnitude limit for spectroscopic redshifts extends beyond $m_{g}\sim 21$ (Khostovan et al., 2025), we consider a conservative apparent magnitude cutoff and remove all FRBs in hosts galaxies with $m_{g}>20.7$. Our results are shown in Figure 5–as expected, the cross-correlation signal is unaffected by the magnitude selection effect on the FRB host galaxy.
3.1.3 Host DM
Measurements of the FRB DM include a component from the FRB’s host galaxy and local environment, typically of the order of tens to a few hundreds of pc cm-3 (Cordes et al., 2022). For studies that are trying to measure the electron density of the IGM or intervening halos (Connor & Ravi, 2022; Connor et al., 2024; Sharma et al., 2025), the DM contribution from the host galaxy is a nuisance term that is typically forward modeled. This is further complicated by the fact that the host DM distributions of FRBs are likely to be redshift dependent Medlock et al. (2025). One advantage of FRB-DM galaxy cross-correlations is that the host DM of the FRB is not expected to correlate with the foreground galaxies, and hence the problem of assuming a functional form for the host DM contribution is avoided.
As we discussed above, the resolution of our electron density bins is insufficiently fine to resolve host DMs. This allows us to verify the assumption that the FRB host galaxy DM contribution to the cross-correlation is negligible by artificially injecting a host DM on top of the DM from the ray tracing. To draw the host DM values, we use the log-normal parameterization from Shin et al. (2023):
$P(\mathrm{DM}_{\text{host}})=\frac{1}{\mathrm{DM}_{\text{host}}}\frac{1}{\sigma\sqrt{2\pi}}\exp\left[-\frac{(\ln\mathrm{DM}_{\text{host}}-\mu)^{2}}{2\sigma^{2}}\right]$ (20)
with best-fit values $\mu=1.93/\log_{10}e$ and $\sigma=0.41/\log_{10}e$. In Figure 6, we compare the cross-correlation measurement without the injected host DM and with the injection. While the host DM contribution increases the variance of the cross-correlation measurement, it does not introduce a bias.
3.2 Propagation selection effects
Now we turn to selection effects on observational properties of FRBs that also arise from propagation effects, namely scattering and DM. Although the host properties of an FRB contribute to both the observed scattering and DM, we ignore those effects here since we have demonstrated that the cross correlation is robust to the properties of the FRB host in Section 3.1.
3.2.1 DM dependent selection effects
To model the DM-dependent selection effects, we use a log-normal probability function based off of the selection function in CHIME/FRB Collaboration et al. (2021b)
$P_{\text{det}}(\mathrm{DM})=\exp\left\{-\frac{2}{3}[\log_{10}(a\cdot\mathrm{DM})-3]^{2}\right\}$ (21)
where DM is in units of $\mathrm{pc}\,\mathrm{cm}^{-3}$ and $a$ is a free scaling parameter. The value $a=1$ roughly corresponds to the DM-dependent selection effect found in CHIME/FRB Collaboration et al. (2021b), while $a=2$ very roughly scales it down to the FRB redshift range in this study, $z_{\text{host}}<0.4$. We compare our results when applying this cut with $a=2$ to the fiducial case in Figure 7, and find that the cross-correlation is largely unaffected.
We also test two additional DM dependent selection functions that are more aggressive: one using Equation 21 with $a=5$, and another as a stepwise function cutting all FRBs with DM$>600$ pc/cm-3. The stepwise function is motivated by the limits on realtime memory buffers that can only hold a fixed amount of data and therefore a limited dispersive sweep while FRBs are searched for in backend systems (e.g. $\sim$ 1000 pc/cm-3 for CHIME (Michilli et al., 2021)).
Our results are shown in Figure 8. We find that a sufficiently aggressive DM dependent selection effect can significantly reduce the amplitude of the cross-correlation signal. Removing the 5% of FRBs with the largest DMs results in the cross-correlation amplitude dropping below $50\%$ at angular scales smaller than $\ell>1000$ (corresponding roughly to physical sizes smaller than $\sim$ 3 Mpc at z=0.35). This implies that the majority of the signal on the scales of galaxy clusters is contained in top 10% of FRB DM contributions, since those trace the largest over-densities.
Figure 9: Topmost panel: histogram of probabilities that an FRB in our simulation is excluded in the cross-correlation due to the scattering selection function presented in Equation 22. Second panel: change in DM distribution of FRB sample after applying scattering selection effect. Bottom two panels: The lowermost panel shows the cross-correlation amplitudes after applying scattering selection effects (lavender) versus the fiducial sample (dotted magenta), with the ratios shown in the middle panel. Scattering from intervening haloes does not significantly affect of the DM-galaxy cross-correlation for our FRB sample.
3.2.2 Scattering
If scattering preferentially removes FRBs with large electron overdensities along the sightline, it is expected to decrease the amplitude of the cross-correlation signal.
We model the probability that an FRB is not detected due to scattering from intervening galaxies with a simple bell curve with a characteristic time-scale of $1\,\mathrm{ms}$:
$P_{\rm{scatter}}(\tau)=1-2^{-\left(\frac{\tau}{1\,\mathrm{ms}}\right)^{2}}$ (22)
We simulate scattering timescales for our FRBs using a simplified version of the cloudlet model based on the work presented by Ocker et al. (2021). We stress that while the purely-geometrical scattering model we choose to adopt here is very simplified, the purpose of this analysis is to demonstrate the extent to which scattering can affect the observed measurement of $C_{\ell}^{Dg}$, and we leave more realistic modeling of scattering (e.g. incorporating gas temperature) in ray tracing simulations to future work.
We calculate the scattering timescale contribution for an FRB observed at frequency $\nu$ with an intervening galaxy halo at redshift $z_{\ell}$ and impact parameter $b_{\perp}$ to be
$\tau=0.3\,\rm{ms}\,\frac{G}{(1+z_{\ell})^{3}(\frac{\nu}{\rm{1\,Ghz}})^{4}}2^{-\left(\frac{b_{\perp}}{4\,\mathrm{kpc}}\right)^{2}}$ (23)
which is based on the simulated curves presented in Figure 2 of Ocker et al. (2021). G, also known as the “geometric leverage factor”, takes the form
$G_{g}=\frac{2d_{fg}d_{go}}{d_{fo}L}$ (24)
where $d_{fg},d_{go},d_{fo}$ are the FRB-galaxy, galaxy-observer, and FRB-observer distances, respectively, and $L$ denotes the thickness of the scattering medium. We take $L$ to be 1 kpc (Cordes et al., 2022). We take our observations to be at $\nu=600$Mhz based on the central observing frequency of CHIME, and the total scattering timescale $\tau$ is given by the sum over all intervening galaxy contributions.
Figure 9 shows the cross-correlation with the scattering model applied. We find that scattering due to intervening halos does not significantly affect the cross-correlation signal. However, scattering from intervening material is expected to increase at higher redshifts due to the geometric leverage $G$ (Ocker et al., 2021). As such, while the cross-correlation appears robust to scattering for our local sample ($z<0.4$), this may not be the case for a catalog of high redshift FRBs, which requires further investigation in future studies.
4 Discussion and Conclusion
In this work, we have investigated the impact of expected observational selection effects on dispersion–galaxy angular cross-correlations using ray tracing simulations with the Illustris TNG300-1 simulation. Our analysis focused on a local sample of 3000 FRBs within the redshift range 0.3 $\leq$ z $\leq$ 0.4 cross-correlated with foreground galaxies in a redshift range of 0.2 $\leq$ z $\leq$ 0.3.
Limitations of our simulation framework prevent us from matching our simulated survey parameters to realistic surveys; we can only simulate patches of the sky 6 degrees in extent and ray trace up to a maxiumum redshift of 0.4. In contrast, CHIME/FRB surveys over half the sky (CHIME/FRB Collaboration et al., 2018) with its FRBs extending to redshift 1 (Shin et al., 2023). As such, our goal is not to precisely quantify the selection effects for any given survey, but to get a first sense of what selection effects are likely to be important. This will provide guidance in planning figure surveys, which will ultimately need to perform simulations matched to their survey to interpret results. Such larger-scale simulations should be feasible in the future using our ray-tracing framework an larger-box simulations such as MilleniumTNG (Pakmor et al., 2023).
[…]
We demonstrated that these cross-correlations are robust to a large swath of potential systematic errors. These include all effects confined to the host galaxy plane, such as variations in FRB host galaxy properties, host DM contributions, and optical follow-up selection effects biased against FRBs with dim galaxy hosts. This result is perhaps expected: the power of cross-correlation studies, in general, are that they are robust to effects confined to only one of the tracers.
We’ve conducted this analysis assuming properties are independent, although in practice this is complicated by the fact that some FRB observables (e.g. host DM and host luminosity) may be correlated in a complicated way. However, we maintain that this can be expected to have a very small effect since observables that are “local” to the FRB are not correlated with the DM overdensities and underdensities within the plane of the galaxies that are used in the cross correlation. That is to say the only FRB selection effects that can bias the power spectrum measurement are DM dependent selection effects and other selection effects on “propagation” observables that have a non-zero covariance with the DM within the foreground galaxy plane (such as scattering).
A priori, selection biases that related to propagation effects are more concerning, since they correlate with the foreground galaxies. Scattering was initially concerning, since it provides a mechanism by which the lines of sight that pass closest to foreground galaxies are lost. However, in our simulations, which employ a simple model for both scattering and the selection bias against it, we observe no effect on the cross-power spectrum. While scattering due to gas external to galactic discs is highly uncertain, within our current understanding it appears not to be a concern for large-scale structure studies at low $(z<0.4)$ redshifts.
As such, of the systematic errors we have considered, selection bias against DM itself is the sole concern for dispersion–galaxy cross-correlations. The magnitude of the bias depends on the details of the selection function. Roughly scaling CHIME’s FRB selection function to the DMs in our simulations, we expect the cross-power to be suppressed by order 10%, worsening on small angular scales. Yet more pernicious is a sharp cutoff in a survey’s ability to detect or localize high DMs, such as is present in CHIME’s ability to capture baseband data, a requirement for VLBI localization with the Outriggers (FRB Collaboration et al., 2025b). We find that excluding the 10% most-dispersed FRBs reduces the amplitude of the cross-correlation signal by a factor of two on angular scales of approximately $0.1^{\circ}$ (corresponding to $\sim$ Mpc scales at our simulated galaxy redshift range).
There are a number of ways that future measurements of the dispersion–galaxy cross-power spectrum could overcome the bias from DM-dependent selection effects. The selection effects could be simulated—using the framework developed here but with survey parameters matched to the measurement—and accounted for in any interpretations. Alternately, it may be possible to develop an analytic model.
Another strategy is to reduce or eliminate the selection effects in the first place. The sharp cutoff bias in CHIME could be overcome via memory upgrades for the baseband systems, including the Outriggers. The FRB detection selection function, coming mainly from dispersive smearing reducing S/N, should be less severe for instruments operating at higher frequencies or with better frequency resolution. Even without instrument changes, a bias-free subsample could be created by selecting detected events above a fixed fluence, which is invariant to dispersion.
Our findings suggest that cross-correlation techniques remain a promising method to probe the distribution of baryons using FRB DMs, but selection effects to dispersion must be properly accounted for to avoid biased measurements. As next-generation radio telescopes such as CHORD, DSA-3000, and BURSTT begin to collect larger and more precise samples of FRBs, our results provide a framework for understanding and mitigating selection biases that may impact FRB surveys aimed at probing the large-scale distribution of baryons.
5 Data Availability
All code, including a FRB ray-tracing package illustris_frb and scripts for processing Illustris data and running experiments, is available on Github [^4] https://github.com/aqcheng/illustris_frb, along with the notebook for generating all data and figures in this paper.
acknowledgments
We thank Kaitlyn Shin for their helpful comments on this paper. We would also like to thank Ryan Raikman, Ralf Konietzka, Liam Connor, and Calvin Leung for conversations that have helped inform this paper.
A.Q.C. is partially supported by the MIT UROP program. K.W.M. holds the Adam J. Burgasser Chair in Astrophysics and received support from NSF grant 2018490.
Appendix A Power spectrum estimator derivation
In this section, we show that the OQE (Equations 12 and 13) reduces to the naive estimator (Equation 9) in the case of complete sky coverage, i.e. when $\mathbf{D}^{-1}$ and $\mathbf{G}^{-1}$ are proportional to the identity. First, let us compute the derivative of the covariance matrix $\mathbf{C}_{,\lambda}$. Recall from Equation 6 that the 2D DFT over a flat square sky patch of $N\times N$ pixels for some field $f_{\vec{\theta}}$ is
$\tilde{f}(\vec{\ell})$ $=\frac{A}{N}\sum_{\vec{\theta}}Q_{\vec{\ell}\vec{\theta}}f_{\vec{\theta}}$ (A1)
$f(\vec{\theta})$ $=\frac{N}{A}\sum_{\vec{\ell}}Q^{\dagger}_{\vec{\theta}\vec{\ell}}\tilde{f}_{\vec{\ell}}$ (A2)
where we have defined the orthonormal operators
$Q_{\vec{\ell}\vec{\theta}}$ $=\frac{1}{N}e^{-i\vec{\ell}\cdot\vec{\theta}}$ (A3)
$Q^{\dagger}_{\vec{\theta}\vec{\ell}}$ $=\frac{1}{N}e^{i\vec{\ell}\cdot\vec{\theta}}$ (A4)
where $\vec{\theta}$ is the position vector in the 2-dimensional configuration space and $\vec{\ell}$ is the wavevector.
Therefore, $\mathbf{C}_{,\lambda}$ can be computed as
$\begin{split}(C_{,\lambda})_{\vec{\theta}\vec{\theta}^{\prime}}=\frac{\partial C_{\vec{\theta}\vec{\theta}^{\prime}}}{\partial C^{Dg}_{\lambda}}&=\frac{\partial\left\langle d_{\vec{\theta}}g^{*}_{\vec{\theta}^{\prime}}\right\rangle}{\partial C^{Dg}_{\lambda}}\\ &=\frac{N^{2}}{A^{2}}\frac{\partial}{\partial C^{Dg}_{\lambda}}\sum_{\vec{\ell},\,\vec{\ell}^{\prime}}Q^{\dagger}_{\vec{\theta}\vec{\ell}}\left\langle\tilde{d}_{\vec{\ell}}\hbox{ }\tilde{g}^{*}_{\vec{\ell}^{\prime}}\right\rangle Q_{\vec{\ell}^{\prime}\vec{\theta}^{\prime}}\\ &=\frac{N^{2}}{A}\frac{\partial}{\partial C^{Dg}_{\lambda}}\sum_{\vec{\ell}}Q^{\dagger}_{\vec{\theta}\vec{\ell}}\hbox{ }C^{Dg}_{\ell}Q_{\vec{\ell}\vec{\theta}^{\prime}}\\ &=\frac{N^{2}}{A}\sum_{\vec{\ell}\in\lambda}Q^{\dagger}_{\vec{\theta}\vec{\ell}}\hbox{ }Q_{\vec{\ell}\vec{\theta}^{\prime}},\end{split}$ (A5)
where we substituted the definition of the cross-correlation Equation 7 on the third line, and substituted $C_{\ell}^{Dg}\approx C_{\lambda}^{Dg}$ for $|\vec{\ell}|\in\lambda$ in the last line. With complete sky coverage, the Equation 12 reduces to
$C_{\lambda}^{Dg}=\sum_{\lambda^{\prime}}\mathrm{Tr}[\mathbf{C}_{,\lambda}\mathbf{C}_{,\lambda^{\prime}}]^{-1}\mathbf{g}^{\dagger}\mathbf{C}_{,\lambda^{\prime}}\mathbf{d}$ (A6)
With the result from Equation A5, we can compute the first term as
$\begin{split}\mathrm{Tr}[\mathbf{C}_{,\lambda}\mathbf{C}_{,\lambda^{\prime}}]&=\frac{N^{4}}{A^{2}}\sum_{|\vec{\ell}|\in\lambda}\sum_{|\vec{\ell}^{\prime}|\in\lambda^{\prime}}\sum_{\vec{\theta},\vec{\theta}^{\prime}}Q_{\vec{\ell}\vec{\theta}}\,Q^{\dagger}_{\vec{\theta}\vec{\ell}^{\prime}}Q_{\vec{\ell}^{\prime}\vec{\theta}^{\prime}}Q^{\dagger}_{\vec{\theta}^{\prime}\vec{\ell}}\\ &=\frac{N^{4}}{A^{2}}\sum_{|\vec{\ell}|\in\lambda}\sum_{|\vec{\ell}^{\prime}|\in\lambda^{\prime}}\delta_{\vec{\ell}\vec{\ell}^{\prime}}=\frac{N^{4}}{A^{2}}N_{\lambda},\end{split}$ (A7)
where $N_{\lambda}$ is the number of modes in the bandpower $\lambda$. The remainder of Equation A6 is computed as
$\begin{split}\mathbf{d}^{\mathrm{T}}\mathbf{C}_{,\lambda^{\prime}}\mathbf{g}&=\frac{N^{2}}{A}\sum_{|\vec{\ell}|\in\lambda}\sum_{\vec{\theta},\,\vec{\theta}^{\prime}}d_{\vec{\theta}}\,Q^{\dagger}_{\vec{\theta}\vec{\ell}}Q_{\vec{\ell}\vec{\theta}^{\prime}}\,g_{\vec{\theta}^{\prime}}\\ &=\frac{N^{4}}{A^{3}}\sum_{|\vec{\ell}|\in\lambda}\tilde{d}^{*}_{\vec{\ell}}\,\tilde{g}_{\vec{\ell}}\end{split}$ (A8)
Thus Equation A6 is equivalent to
$\hat{C}_{\lambda}^{Dg}=\frac{1}{A\,N_{\lambda}}\sum_{|\vec{\ell}|\in\lambda}\tilde{d}^{*}_{\vec{\ell}}\,\tilde{g}_{\vec{\ell}},$ (A9)
exactly the naive estimator in Equation 9.
Finally, we show that the OQE in our implementation is on average equivalent to the so-called pseudo-$C_{\ell}$ estimator (Alonso et al., 2019), which is the naive estimator scaled by the available sky fraction:
$\langle\hat{C}_{\lambda}^{Dg}\rangle=\frac{1}{f_{sky}}\frac{1}{AN_{\lambda}}\sum_{\vec{\ell}\in\lambda}\left\langle\left(\tilde{d}^{m}_{\vec{\ell}}\right)^{*}\tilde{g}^{m}_{\vec{\ell}}\right\rangle,$ (A10)
With a general window function, the term $\mathbf{d}^{\mathrm{T}}\mathbf{C}_{,\lambda^{\prime}}\mathbf{g}$ stays the same since $\mathbf{D}^{-1}\mathbf{d}^{m}=\mathbf{d}^{m}$ and $\mathbf{G}^{-1}\mathbf{g}^{m}=\mathbf{g}^{m}$, i.e. applying a mask to a masked field does not change it. The trace term Equation A7, however, is generalized to
$\begin{split}\mathrm{Tr}[\mathbf{C}_{,\lambda}\mathbf{G}^{-1}\mathbf{C}_{,\lambda^{\prime}}\mathbf{D}^{-1}]&=\frac{N^{4}}{A^{2}}\sum_{|\vec{\ell}|\in\lambda}\sum_{|\vec{\ell}^{\prime}|\in\lambda^{\prime}}\sum_{\vec{\theta},\vec{\theta}^{\prime}}Q^{\dagger}_{\vec{\theta}\vec{\ell}}~{}Q_{\vec{\ell}\vec{\theta}^{\prime}}\,(G^{-1})_{\vec{\theta}^{\prime}}Q^{\dagger}_{\vec{\theta}^{\prime}\vec{\ell}^{\prime}}~{}Q_{\vec{\ell}^{\prime}\vec{\theta}}(D^{-1})_{\vec{\theta}}\\ &=\frac{N^{4}}{A^{2}}\sum_{|\vec{\ell}|\in\lambda}\sum_{|\vec{\ell}^{\prime}|\in\lambda^{\prime}}(\mathbf{Q}\mathbf{D}^{-1}\mathbf{Q}^{\dagger})_{\vec{\ell}\vec{\ell}^{\prime}}(\mathbf{Q}\mathbf{G}^{-1}\mathbf{Q}^{\dagger})_{\vec{\ell}^{\prime}\vec{\ell}}.\end{split}$ (A11)
In our case, we have complete sky coverage for galaxies ($\mathbf{G}^{-1}=\mathbf{I}$), while FRBs occupy a random fraction $f_{sky}$ of pixels in the sky region. Therefore, $\langle\mathbf{G}^{-1}\rangle=f_{sky}\mathbf{I}$ and
$\langle\mathrm{Tr}[\mathbf{C}_{,\lambda}\mathbf{G}^{-1}\mathbf{C}_{,\lambda^{\prime}}\mathbf{D}^{-1}]\rangle=\frac{N^{4}}{A^{2}}N_{\lambda}f_{sky}.$ (A12)
Taking the expectation value of $C_{\lambda}^{Dg}$ in Equation 12 and assuming that the survey window is uncorrelated with the DM field then yields the result in Equation A10.
Appendix B Simulation sky patch selection
We consider two conditions in our selection of good sky patches: (1) no ray within the region will intersect itself, and (2) no two rays within the region will intersect the same patch of the simulation box.
The first problem of the self-intersecting ray is manifest in any field with periodic boundary conditions. For example, naively following a ray parallel to any axis would result in repeating the same exact path through the simulation box, once we enforce periodic boundary conditions. A ray to a FRB at $z=0.4$ for a simulation box of size 300 cMpc would cross the same structures $5$ times.
To identify which sky directions are problematic up to $z=0.4$, recall that periodic boundary conditions identifies any point $\vec{x}$ within the simulation box with points $\vec{x}+b(k,m,n)$, where $b$ is the simulation box size and $k,m,n$ are integers. Now, consider the vector corresponding to the ray, $\vec{v}=\vec{x}_{FRB}-\vec{x}_{0}$. Condition (1) is equivalent to the condition that $\vec{x}_{0}+b(k,m,n)$ does not lie anywhere along the ray for all $k,m,n$. Therefore, we can find all problematic sky directions by computing all $|b(k,m,n)|<|\vec{v}|=\chi(z=0.4)$, where $\chi$ is the comoving distance. Because our simulation box is gridded into bins of size $500\,\mathrm{ckpc}/h$, each of these sky directions excludes a circular sky area of radius $\theta(z)=\frac{500/(1+z)\,\mathrm{kpc}/h}{D_{A}(z)}$ evaluated at $z=0.4$, where $D_{A}$ is the angular diameter distance. In a flat cosmology, this is a radius of $\theta(z)=\frac{500\,\mathrm{ckpc}/h}{\chi(z)}\approx 5\cdot 10^{-4}\,\mathrm{rad}$ at $z=0.4$.
The second problem is the intersection of nearby sightlines, e.g. a grid of FRBs within a small sky patch. This can occur when stacking periodic boxes; see Figure 10. In order to avoid this, the sightline $\vec{v}$ must be sufficiently far apart from its images $\vec{v}+b(k,m,n)$, in particular its adjacent images (i.e. $k,m,n\in\{0,1\})$. An “image” here refers to any identical path to the sightline as enforced by the periodic boundary condition. Figure 10 visualizes this issue: sightline A (solid blue) must be sufficiently far apart from its images (dashed blue) such that the other sightlines in the same sky region do not intersect them.
[…]
Figure 10: Parallel sightlines can intersect each other in different periodic boxes. Here, nearby sightlines B and C (red) intersect the images (dashed blue lines) of sightline A in adjacent boxes. Notably, a large intervening structure (e.g. a galaxy cluster) can yield spurious spatial patterns, as sightlines A, B, and C would have DM contributions from the same structure (oval), whereas sightlines in between them would not.
Let $\vec{u}$ be any $b(k,m,n)$ with $k,m,n\in\mathbb{Z}$. The smallest distance between two rays $\vec{v}$ and $\vec{v}+\vec{u}$ is given by $\hat{v}\times\vec{u}$, where $\hat{v}=\frac{\vec{v}}{|\vec{v}|}$ is the unit vector corresponding to $\vec{v}$ and uniquely identifies a sky direction. Thus for any given sky direction $\hat{v}$, we can compute the distance to the closest adjacent image
$s(\hat{v})=\min_{\vec{u}}\left|\hat{v}\times\vec{u}\right|$ (B1)
Therefore, for a given redshift $z$ at which the FRBs will be placed, the selection of a “good” sky patch of angular size $\theta$ is the requirement
$s(\hat{v})>\theta\,\chi(z).$ (B2)
Applying both of these conditions to yields the map of good sightlines, as shown by the white regions in Figure 1 for $\theta=0.1\,\mathrm{rad}$ and $z=0.4$. The requirement (2) that nearby rays cannot intersect in adjacent boxes is extremely restrictive. Furthermore, the “good” sky areas shrink when we ray trace to greater distances, as the ray crosses more periodic boxes. This restriction sets a limit for the sky region sizes and FRB redshifts we can explore within IllustrisTNG. It is for this reason that we choose a sky patch size of $0.1$ rad and a maximum redshift of $z=0.4$.